# HYDRA: Unifying Multi-modal Generation and Understanding via Representation-Harmonized Tokenization

### 16 Mar 2026 v1

@@ F.1 Multi-modal Understanding Benchmarks
To comprehensively evaluate the perception and reasoning capabilities of HYDRA, we employ eight diverse benchmarks covering general understanding, expert knowledge, and fine-grained visual perception. We utilize MME (mme) to measure comprehensive perception and cognition, reporting MME-S as the aggregate sum of both scores. MMBench (MMB) (mmbench) is employed for its robust circular evaluation strategy, assessing fine-grained perception and logical reasoning. To ensure rational visual dependency, we evaluate on MMStar (chen2024mmstar), a curated benchmark that filters out samples solvable by text alone. Additionally, SEED-Bench (SEED) (li2023seed) serves as a large-scale testbed for generative and discriminative comprehension across spatial and temporal dimensions. For high-level discipline-specific reasoning, we employ MMMU (yue2024mmmu), a massive multi-discipline benchmark demanding expert-level reasoning across fields such as art, science, and engineering. This is complemented by AI2D (kembhavi2016ai2d), a specialized dataset focused on understanding and answering questions about scientific diagrams and textbook illustrations. To assess text-centric visual understanding, we use OCRBench (liu2024ocrbench), a dedicated benchmark for optical character recognition tasks. Furthermore, RealWorldQA (RWQA)  is utilized to evaluate spatial reasoning and physical understanding in diverse real-world environments.
(a) AI2D (kembhavi2016ai2d).
(b) MMMU (yue2024mmmu).
(c) OCRBench (liu2024ocrbench).
(d) MMBench (mmbench).
(e) MMStar (chen2024mmstar).
(f) SEED-Bench (li2023seed).
@@ F.2 CKNNA
CKNNA (Centered Kernel Nearest-Neighbor Alignment) is a similarity metric designed to compare representations by emphasizing local relational structure rather than global correspondence. Unlike Centered Kernel Alignment (CKA; kornblith2019similarity), which aggregates correlations over all sample pairs, CKNNA restricts the comparison to pairs that are mutually close in both representation spaces, thereby providing a more flexible and locality-aware measure. Our formulation follows the core principles introduced in (huh2024platonic; repa), while adopting an equivalent but implementation-aligned presentation.
@@
In practice, we uniformly sample 10,00010{,}000 images from the ImageNet validation set (imagenet) and compute CKNNA with $k=10$. As observed in (huh2024platonic), restricting the comparison to small neighborhoods yields a more sensitive measure of representational agreement.
@@ G.1 Image Reconstruction
Figure 10: Qualitative comparison of image reconstruction. We compare visual results from our HYDRA between RAE (rae), SD1.5 (sdxl), VTP (yao2025vtp), and FLUX (flux).
@@ G.2 Image Generation
Figure 11: Qualitative comparison of text-to-image generation. We compare visual generation between Harmon (wu2025harmonizing), Show-o2 (xie2025show), Bagel (bagel), Janus-pro (wu2025janus), and Blip-3o (blip3-o).
Figure 12: Qualitative comparison of text-to-image generation. We compare visual generation results from our HYDRA between Harmon (wu2025harmonizing), Show-o2 (xie2025show), Bagel (bagel), Janus-pro (wu2025janus), and Blip-3o (blip3-o).

### 17 Mar 2026 v2

@@ F.1 Multi-modal Understanding Benchmarks
To comprehensively evaluate the perception and reasoning capabilities of HYDRA, we employ eight diverse benchmarks covering general understanding, expert knowledge, and fine-grained visual perception. We utilize MME (Fu et al., 2023) to measure comprehensive perception and cognition, reporting MME-S as the aggregate sum of both scores. MMBench (MMB) (Liu et al., 2024a) is employed for its robust circular evaluation strategy, assessing fine-grained perception and logical reasoning. To ensure rational visual dependency, we evaluate on MMStar (Chen et al., 2024a), a curated benchmark that filters out samples solvable by text alone. Additionally, SEED-Bench (SEED) (Li et al., 2023a) serves as a large-scale testbed for generative and discriminative comprehension across spatial and temporal dimensions. For high-level discipline-specific reasoning, we employ MMMU (Yue et al., 2024), a massive multi-discipline benchmark demanding expert-level reasoning across fields such as art, science, and engineering. This is complemented by AI2D (Kembhavi et al., 2016), a specialized dataset focused on understanding and answering questions about scientific diagrams and textbook illustrations. To assess text-centric visual understanding, we use OCRBench (Liu et al., 2024b), a dedicated benchmark for optical character recognition tasks. Furthermore, RealWorldQA (RWQA)  is utilized to evaluate spatial reasoning and physical understanding in diverse real-world environments.
(a) AI2D (Kembhavi et al., 2016).
(b) MMMU (Yue et al., 2024).
(c) OCRBench (Liu et al., 2024b).
(d) MMBench (Liu et al., 2024a).
(e) MMStar (Chen et al., 2024a).
(f) SEED-Bench (Li et al., 2023a).
@@ F.2 CKNNA
CKNNA (Centered Kernel Nearest-Neighbor Alignment) is a similarity metric designed to compare representations by emphasizing local relational structure rather than global correspondence. Unlike Centered Kernel Alignment (CKA; Kornblith et al. 2019), which aggregates correlations over all sample pairs, CKNNA restricts the comparison to pairs that are mutually close in both representation spaces, thereby providing a more flexible and locality-aware measure. Our formulation follows the core principles introduced in (Huh et al., 2024; Yu et al., 2024), while adopting an equivalent but implementation-aligned presentation.
@@
In practice, we uniformly sample 10,00010{,}000 images from the ImageNet validation set (Russakovsky et al., 2014) and compute CKNNA with $k=10$. As observed in (Huh et al., 2024), restricting the comparison to small neighborhoods yields a more sensitive measure of representational agreement.
@@ G.1 Image Reconstruction
Figure 10: Qualitative comparison of image reconstruction. We compare visual results from our HYDRA between RAE (Zheng et al., 2025), SD1.5 (Podell et al., 2023), VTP (Yao et al., 2025a), and FLUX (Labs et al., 2025).
@@ G.2 Image Generation
Figure 11: Qualitative comparison of text-to-image generation. We compare visual generation between Harmon (Wu et al., 2025d), Show-o2 (Xie et al., 2025a), Bagel (Deng et al., 2025a), Janus-pro (Wu et al., 2025b), and Blip-3o (Chen et al., 2025a).
Figure 12: Qualitative comparison of text-to-image generation. We compare visual generation results from our HYDRA between Harmon (Wu et al., 2025d), Show-o2 (Xie et al., 2025a), Bagel (Deng et al., 2025a), Janus-pro (Wu et al., 2025b), and Blip-3o (Chen et al., 2025a).